An aerostat of mass M starts coming down with a constant acceleration W. Determine the ballest mass to be dumped for the aerostat to reach the upward acceleration of the same magnitude. The air drag is to be neglected.

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

there are two forces acting on the aerostat; gravity and vertically upward 'f', MW=Mg-f, f=M(g-w),let ballast mass to be dumped be Md

f-(M-Mdg)=W(M-Md), solve the equation to obtain the answer.

 

ADITYA SONI
10 years ago

IT IS A QUIESTION OF I.E. IRODOV

SOLUTION:

in the beginning we see that

#mg-f=wm (eqn 1)

let us say that the dumped weight be δm

#after dumping we say that

f-(m-δm)=w(m-δ)  (eqn2)

we got two eqn

now what are u waiting for solve it